# Hyperparameter Optimization in Machine Learning

### v1

@@ 8.1.1 Multi-objective optimization
Early efforts on HPO focused on designing evolutionary optimization techniques for neural networks and SVMs [Yao1999:Evolving, Nolfi2002:Evolution, Igel2005:Evolutionary, friedrichs2005evolutionary]. Evolutionary techniques have also been applied to multi-objective HPO problems [Jin2006:Multi]. To mitigate the large computational cost of evaluating a single hyperparameter configuration (e.g., training a DNN), sample-efficient Bayesian optimization (BO) techniques have been introduced [hutter_sequential_2011, snoek_practical_2012, bergstra2011algorithms].
Multi-objective Bayesian optimization (MBO) techniques can be categorized into three classes. First, scalarization-based MBO methods map the vector of all objectives to a scalar [Knowles2006:Parego, Zhang2009:Expensive, Nakayama2009:Sequential, Paria2019:Flexible, Golovin2020:Random] and then use conventional single-objective (SO) BO techniques. The scalarization is defined by using a weight vector $\mathbf{w}\in\mathbb{R}^{n}$, solving the MO minimization problem consists of solving the following optimization problem:
$\min_{\lambda\in\Lambda}V(f(\lambda),\mathbf{w})$ (8.1)
where $V:\mathbb{R}^{n}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a scalarization function. While these methods are comparatively easy to implement and scale gracefully with the number of objectives, they do not utilize information about the overall geometry of the current Pareto front approximation. Some common choices for $V$ are:
@@
  • the scalarization used in [Knowles2006:Parego], called ParEGO: $V_{\text{ParEGO}}(f(\lambda),\mathbf{w})=\max_{j\in\{1,\dots,n\}}(w_{j}f(\lambda)_{j})+\rho(f(\lambda)^{\top}\mathbf{w})$; the scalarization used in [Golovin2020]: $V_{\text{Golovin}}(f(\lambda),\mathbf{w})=\min_{j\in\{1,\dots,n\}}(\max(0,f(\lambda)_{j}/w_{j}))^{n}$.
@@
Finally, information-theoretic MBO approaches aim to select points that reduce uncertainty about the location of the Pareto front. Methods in this class tend to be more sample efficient and scale better with the number of objectives. PAL [Zuluaga2013:Active] iteratively reduces the size of a discrete uncertainty set. PESMO [Hernandez2016:Predictive] adapts the predictive entropy search (PES) [Hernandez2014:Predictive] criterion. Pareto-frontier entropy search (PFES) [Suzuki2019:Multi] is suitable when dealing with decoupled objectives. MESMO [Belakaria2019:Max] builds on the max-value entropy-search criterion [Wang2017:Max] and enjoys an asymptotic regret bound. Building on MESMO, two very recent works have proposed MF-OSEMO [Belakaria2020:Multi] and iMOCA [Belakaria2020:Information], two multi-fidelity based information-theoretic MBO techniques which internally use multi-fidelity Gaussian processes.
@@ 8.1.2 Constrained hyperparameter optimization
Our goal in a constrained optimization scenario is to minimize the target black-box $f(\lambda)$, subject to a constraint $c(\lambda)\leq\delta$. Here, we limit our attention to modeling the feasible region by a single function $c(x)$. As the latent constraints are pairwise conditionally independent, an extension to multiple constraints is straightforward, yet notationally cumbersome. Both $f(\lambda):\Lambda\rightarrow\mathbb{R}$ and $c(x):\Lambda\rightarrow\mathbb{R}$ are unknown and need to be queried sequentially. The constrained optimization problem is defined as follows:
$f_{\star}=\min_{\lambda\in\Lambda}\left\{f(\lambda)\;\|\;c(\lambda)\leq\delta\right\},$ (8.2)
@@
We consider two different setups, depending on what information is observed about the constraint. Many existing works [Gardner14, Gelbart14, Lobato15a] assumes that real-valued feedback is obtained on $c(\lambda)$, just as for $f(\lambda)$. In this case, both latent function can be represented as GPs with Gaussian noise. Unfortunately, this setup does not cover practically important use cases of constrained hyperparameter optimization. For example, if training a DNN fails with an out-of-memory error, we cannot observe the amount of memory requested just before the crash, neither do we usually know the exact amount of memory available on the compute instance. Covering such use cases requires handling binary feedback on $c(x)$, even though this is technically more difficult. We can assume an evaluation returns $z_{f}\sim\mathcal{N}(z_{f}|f(\lambda),\alpha_{f}^{-1})$ and $z_{c}\in\{-1,+1\}$, where $z_{c}=-1$ for a feasible, $z_{c}=+1$ for an unfeasible point. Here, $z_{f}$ is the observed value of the objective function, while $z_{c}$ is the observed value of the constraint. We never observe the latent constraint function $c(\lambda)$ directly. We assume $z_{c}\sim\sigma(z_{c}c(\lambda))$, where $\sigma(t)=\frac{1}{1+e^{-t}}$ is the logistic sigmoid, but other choices are possible. We can then rewrite the constrained optimization problem (8.2) as follows:
$y_{\star}=\min_{x{}\in\mathcal{X}}\left\{y(x{})\;\|\;P(z_{c}=+1|x{})=\sigma(c(x{}))\leq\sigma(\delta)\right\}.$
@@
The most established technique to tackle constrained BO is constrained EI (cEI) [Gardner14, Gelbart14, snoek_scalable_2015, letham2019]. If the constraint is denoted by $c(\lambda)\leq\delta$, a separate regression model is used to learn the constraint function $c(\lambda)$ (typically a GP), and EI is modified in two ways. First, the expected amount of improvement of an evaluation is computed only with respect to the current feasible minimum. Second, hyperparameters with a large probability of satisfying the constraint are encouraged by optimizing $cEI(\lambda)=P\{c(\lambda)\leq\delta\}EI(\lambda)$, where $P\{c(\lambda)\leq\delta\}$ is the posterior probability of $\lambda$ being feasible under the constraint model, and $EI(\lambda)$ is the standard EI acquisition function, see Section 4.2.
Several issues with cEI are detailed in [Lobato15a]. First, the current feasible minimum has to be known, which is problematic if all initial evaluations are unfeasible. A workaround is to use a different acquisition function, initially focused on finding a feasible point [Gelbart14]. In addition, the probability of constraint violation is not explicitly accounted for in cEI. A confidence parameter equivalent to ours features in [Gelbart14], but is only used to recommend the final hyperparameter configuration. Another approach was proposed in [Lobato15a], where PES is extended to the constrained case. Constrained PES (cPES) can outperform cEI and does not require the workarounds mentioned above. However, it is complex to implement, expensive to evaluate, and unsuitable for binary constraint feedback. Drawing from numerical optimization, different generalizations of EI to the constrained case are developed in [Picheny2016] and [Ariafar2019]. The authors represent the constrained minimum by way of Lagrange multipliers, and the resulting query selection problem is solved as a sequence of unconstrained problems. A better option is the extension of another acquisition function to the constrained case, namely Max-value Entropy Search [perrone_cmes]. This acquisition function handles binary constraint feedback, does not require numerical quadrature [Picheny2016], and does not come with a large set of extra hyperparameters [Ariafar2019].
@@ 8.2 Hyperparameter schedules and online HPO
The vast majority of learning algorithms do not produce an hypothesis in one shot, but rather refine it incrementally over the course of many iterations. In the previous chapter we already discussed algorithms designed around this observation, such as those leveraging multi-fidelity approaches (Chapter 5), and hypergradients (Section 7.3), as well as some evolutionary algorithms like population-based training (Section 6.1.2). When this is the case, it can make sense to consider hyperparameter schedules rather than fixed values. One of the most prominent example in the literature are the ubiquitous learning rate schedules when training neural nets [bengio2012practical, loshchilov2017sgdr, goyal2017accurate]. In this section we overview such setup and briefly discuss how one may tackle the HPO problem as an instance of online optimization.
Mimicking the notation of Section 7.3, let us consider an algorithm $A:\mathscr{D}\times\Lambda\to\mathcal{H}$ as a composition of iterates $\Phi_{t}:\mathcal{H}\times\Lambda\times\mathscr{D}\rightarrow\mathcal{H}$. Such transitions maps need not be necessarily differentiable, although several methods that exploit the iterative view rely on this assumption. This time, we let the hyperparameter configuration also vary as a function of the step. That is, given an initial model $h_{0}\in\mathcal{H}$:
$A(\mathcal{D},\lambda)=\Phi_{T}(\Phi_{T-1}(\circ\dots\circ\Phi_{1}(h_{0},\lambda_{1},\mathcal{D}),\lambda_{T-1},\mathcal{D}),\lambda_{T},\mathcal{D}).$ (8.3)
for some $T>0$, where we generally view $\lambda\in\Lambda^{T}$ as an hyperparameter sequence. The iterations represented by the iterates $\Phi_{t}$ can correspond to different granularities or spans of operations within the training process, such as single stochastic gradient descent steps, entire epochs, or any other predefined sequence of operations..
@@
However, perhaps one main advantage of explicitly taking an iterative view is that it enable recasting the HPO problem as an instance of online optimization [shalev2012online], albeit a non-standard one. Online HPO techniques attempt to speed up the HPO procedure by finding a performing hyperparamter schedule in one go, during a single execution of the learning algorithm. A key difficulty is that one does not have access to the true objective $\mathscr{M}$. Indeed, being able to compute (an unbiased estimate of) $\mathscr{M}$ requires terminating the learning algorithm, which would defeat the purpose of online HPO. In this sense, these techniques do not – or rather, cannot – strictly solve Problem (2.4)-(2.5), but rather rely on approximations, heuristics or surrogate objectives. This is akin to what we discussed in Chapter 5, however multi-fidelity techniques only use early observations to take decisions or fit models of the response surface, and do not change hyperparameters during execution. Examples of online HPO include [orabona2016coin, baydin_2017_online, luketina2016scalable, jaderberg2017population, lorenzo2017particle, mackay2019self, donini2019opt, vicol2021unbiased, micaelli2021gradient, chandra2022gradient], with gradient-based techniques, and especially forward iterative differentiation, being proposed by various authors. Online HPO remains an area of active research.
@@ 8.3.1 Search Space
Instead of assigning operations to each node, liu2018darts assigns operations to every edge of the graph. Nodes that accumulate the feature vectors from all incoming edges. Their proposed method DARTS then uses gradient descent to select an operation for each edge.
@@ 8.3.2 Search strategies
In one of the first works on NAS, zoph2016neural used a controller based on a recurrent neural network to suggest new architectures. In each iteration, a set of candidate architectures is sampled and evaluated to collect observations, which are then used to update the controller using reinforcement learning. real2019regularized proposed an evolutionary approach that maintains a population of architectures. In each iteration, a random architecture is drawn from the population and mutated along a single dimension. While these approaches led to architectures outperforming state-of-the-art models at that point, the original methods required substantial computational resources since each network was trained from scratch.
@@ 8.4 Meta-learning
We discuss here meta-learning since, from a certain point of view, it shares some interesting similarities to hyperparameter optimization. This section only focuses on this point of view, in the case of supervised learning. For a more general overview of meta learning, see e.g., [hospedales:2021:metalearning_neural_networks] and references therein.
@@ 8.4.1 Definition
Meta-learning has its own specificity when compared to other settings dealing with several tasks. Compared to multi-task learning [Caruana94, Baxter2000, ruder:2017:overview_multitask_learning], the set of tasks is not fixed in advance. Compared to transfer learning or domain adaptation [pan2009survey, torrey2010transfer, kouw2018introduction], there are several source tasks. Typically, we are interested in the scenario where a very small dataset $\mathcal{D}\leavevmode\nobreak\ \sim p$ will be available for the new task (few-shot learning, e.g. $|\mathcal{D}|=5$), or even a single example (one-shot learning). The main insight is that when the tasks are related, as capture by the distribution $\mathcal{P}$, meta-learning reduces the sample complexity of learning new tasks drawn by the same distribution [Baxter2000, maurer2016benefit].
8.4.2 Meta-learning as bilevel programming
As in [Baxter2000], let us introduce the family of hypothesis spaces $\mathbb{H}^{j}=\{\mathcal{H}^{j}\}$ where each $\mathcal{H}^{j}\in\mathbb{H}^{j}$ is a set of functions $h^{j}_{w,\lambda}:\mathcal{X}\mapsto\mathcal{Y}^{j}$ of the form $h^{j}_{w,\lambda}=\gamma_{w}^{j}\circ\Phi_{\lambda}$, i.e., obtained by composing a shared feature learning function $\Phi_{\lambda}:\mathcal{X}\mapsto\mathcal{F}$, parameterized by $\lambda$ (also called meta-parameter), with a task-specific function $\gamma_{w}^{j}:\mathcal{F}\mapsto\mathcal{Y}_{j}$, parameterized by $w$. The set $\mathcal{F}$ is a suitably chosen feature space, for example $\mathcal{F}=\mathbb{R}^{d}$. This type of function composition is common in several deep learning architectures where $\Phi$ is in charge of representation learning while $\gamma$ could be a classification or regression head. The multi-task network introduced in [Caruana94] and the feature learning neural network described in [Baxter2000] are structured in that way, although all the weights $\lambda$ and $w$ are optimized jointly, unlike the problem formulated below. Note that we have deliberately used the notation $\lambda$ for the parameters shared across the tasks, to highlight the fact that they play a role similar to hyperparameters. In order to construct a “response function,” for every source task $\mathcal{T}^{j}$, $j=1,\dots N$, we split the given dataset $\mathcal{D}^{j}$ into a training set $\mathcal{D}_{\text{tr}}^{j}$ and a validation set $\mathcal{D}_{\text{val}}^{j}$ and define

### v2

@@ 8.1.1 Multi-objective optimization
Initial efforts towards MO optimization for HPO [Jin2006:Multi] were inspired by evolutionary techniques [Yao1999:Evolving, Nolfi2002:Evolution, Igel2005:Evolutionary, friedrichs2005evolutionary]. To mitigate the large computational cost of evaluating a single hyperparameter configuration (e.g., training a DNN), sample-efficient Bayesian optimization (BO) techniques have been introduced [hutter_sequential_2011, snoek_practical_2012, bergstra2011algorithms].
Multi-objective Bayesian optimization (MBO) techniques can be categorized into three classes. First, scalarization-based MBO methods map the vector of all objectives to a scalar [Knowles2006:Parego, Zhang2009:Expensive, Nakayama2009:Sequential, Paria2019:Flexible, Golovin2020:Random] and then use conventional single-objective BO techniques. The scalarization is defined by using a weight vector $\mathbf{w}\in\mathbb{R}^{n}$.
Some common choices for $V$ are:
@@
  • The scalarization used in [Knowles2006:Parego], called ParEGO:
    $V_{\text{ParEGO}}(f(\lambda),\mathbf{w})=\max_{j\in\{1,\dots,n\}}(w_{j}f(\lambda)_{j})+\rho(f(\lambda)^{\top}\mathbf{w});$
  • The scalarization used in [Golovin2020]:
    $V_{\text{Golovin}}(f(\lambda),\mathbf{w})=\min_{j\in\{1,\dots,n\}}(\max(0,f(\lambda)_{j}/w_{j}))^{n}.$
@@
Finally, information-theoretic MBO approaches aim to select points that reduce uncertainty about the location of the Pareto front. Methods in this class tend to be more sample efficient and scale better with the number of objectives. PAL [Zuluaga2013:Active] iteratively reduces the size of a discrete uncertainty set. PESMO [Hernandez2016:Predictive] adapts the predictive entropy search (PES) [Hernandez2014:Predictive] criterion. Pareto-frontier entropy search (PFES) [Suzuki2019:Multi] is suitable when dealing with decoupled objectives. MESMO [Belakaria2019:Max] builds on the max-value entropy-search criterion [Wang2017:Max] and enjoys an asymptotic regret bound. Building on MESMO, two recent works have proposed MF-OSEMO [Belakaria2020:Multi] and iMOCA [Belakaria2020:Information], two multi-fidelity based information-theoretic MBO techniques which internally use multi-fidelity Gaussian processes.
@@ 8.1.2 Constrained hyperparameter optimization
When optimizing for multiple targets, some objectives may have predefined constraints on their values. For example, one might seek to optimize a given objective function while enforcing a strict memory constraint. In this section, we focus on this scenario.
Formally, the goal in a constrained optimization scenario is to minimize the target response function $f(\lambda)$, subject to a constraint $c(\lambda)\leq\delta$. Here, we limit our attention to modeling the feasible region by a single function $c(x)$. As the latent constraints are pairwise conditionally independent, an extension to multiple constraints is straightforward, yet notationally cumbersome. Both $f(\lambda):\Lambda\rightarrow\mathbb{R}$ and $c(x):\Lambda\rightarrow\mathbb{R}$ are unknown and need to be queried sequentially. The constrained optimization problem is defined as follows:
$\min_{\{\lambda\in\Lambda\,:\,c(\lambda)\leq\delta\}}f(\lambda),$ (8.1)
@@
We consider two different setups, depending on what information is observed about the constraint. Many existing works [Gardner14, Gelbart14, Lobato15a] assumes that real-valued feedback is obtained on $c(\lambda)$, just as for $f(\lambda)$. In this case, both latent function can be represented as GPs with Gaussian noise. Unfortunately, this setup does not cover practically important use cases of constrained hyperparameter optimization. For example, if training a DNN fails with an out-of-memory error, we cannot observe the amount of memory requested just before the crash. Rather, we only know that an out-of-memory error occurred, which is a binary outcome. Covering such use cases requires handling binary feedback on $c(x)$, even though this is technically more difficult. We can assume an evaluation returns $z_{f}\sim\mathcal{N}(z_{f}|f(\lambda),\alpha_{f}^{-1})$ and $z_{c}\in\{-1,+1\}$, where $z_{c}=-1$ for a feasible, $z_{c}=+1$ for an unfeasible point. Here, $z_{f}$ is the observed value of the objective function, while $z_{c}$ is the observed value of the constraint. We never observe the latent constraint function $c(\lambda)$ directly. We assume $z_{c}\sim\sigma(z_{c}c(\lambda))$, where $\sigma(t)=\frac{1}{1+e^{-t}}$ is the logistic sigmoid, but other choices are possible. We can then rewrite the constrained optimization problem (8.1) as follows:
$y_{\star}=\min_{\lambda\in\Lambda}\left\{y(\lambda)\;\|\;P(z_{c}=+1|\lambda)=\sigma(c(\lambda))\leq\sigma(\delta)\right\}.$
@@
The most established technique to tackle constrained BO is constrained EI (cEI) [Gardner14, Gelbart14, snoek_scalable_2015, letham2019]. A separate regression model is used to learn the constraint function $c(\lambda)$ (typically a GP), and EI is modified in two ways. First, the expected amount of improvement of an evaluation is computed only with respect to the current feasible minimum. Second, hyperparameters with a large probability of satisfying the constraint are encouraged by optimizing $cEI(\lambda)=\mathbb{P}\{c(\lambda)\leq\delta\}EI(\lambda)$, where $\mathbb{P}\{c(\lambda)\leq\delta\}$ is the posterior probability of $\lambda$ being feasible under the constraint model, and $EI(\lambda)$ is the standard EI acquisition function, see Section 4.2.3.
Several issues with cEI are detailed in [Lobato15a]. First, the current feasible minimum has to be known, which is problematic if all initial evaluations are unfeasible. A workaround is to use a different acquisition function, initially focused on finding a feasible point [Gelbart14]. In addition, the probability of constraint violation is not explicitly accounted for in cEI. A confidence parameter features in [Gelbart14], but is only used to recommend the final hyperparameter configuration. Another approach was proposed in [Lobato15a], where PES is extended to the constrained case. Constrained PES (cPES) can outperform cEI and does not require the workarounds mentioned above. However, it is complex to implement, expensive to evaluate, and unsuitable for binary constraint feedback. Drawing from numerical optimization, different generalizations of EI to the constrained case are developed in [Picheny2016] and [Ariafar2019]. The authors represent the constrained minimum by way of Lagrange multipliers, and the resulting query selection problem is solved as a sequence of unconstrained problems. An alternative option that does not require expensive computations is the extension of another acquisition function to the constrained case, namely Max-value Entropy Search [perrone_cmes]. This acquisition function handles binary constraint feedback, does not require numerical quadrature [Picheny2016], and does not come with a large set of extra hyperparameters [Ariafar2019].
@@ 8.2 Hyperparameter schedules and online HPO
The vast majority of learning algorithms do not produce an hypothesis in one shot, but rather refine it incrementally over the course of many iterations. In the previous chapters we already discussed algorithms designed around this observation, such as those leveraging multi-fidelity approaches (Chapter 5), and hypergradients (Section 7.3), as well as some evolutionary algorithms like population-based training (Section 6.3). When this is the case, it may be sensible to consider hyperparameter schedules rather than fixed values. One of the most prominent example in the literature are the ubiquitous learning rate schedules when training neural nets [bengio2012practical, loshchilov2017sgdr, goyal2017accurate]. In this section we overview such setup and briefly discuss how one may tackle the HPO problem as an instance of online optimization.
Mimicking the notation of Section 7.3, let us consider an algorithm $A:\Lambda\to\mathcal{H}$ as a composition of iterates $\Phi_{t}:\mathcal{H}\times\Lambda\times\rightarrow\mathcal{H}$. 111 We disregard the dependence of the maps on data, for notational simplicity. Such transitions maps need not be necessarily differentiable, although several methods that exploit the iterative view rely on this assumption. This time, we let the hyperparameter configuration also vary as a function of the step. That is, given an initial model $h_{0}\in\mathcal{H}$:
$A(\lambda)=\Phi_{T}(\cdot,\lambda_{T})\circ\Phi_{T-1}(\cdot,\lambda_{T-1})\circ\dots\circ\Phi_{1}(h_{0},\lambda_{1}).$ (8.2)
for some $T>0$, where we generally view $\lambda\in\Lambda^{T}$ as an hyperparameter sequence. The iterations represented by the iterates $\Phi_{t}$ can correspond to different granularities or spans of operations within the training process, such as single stochastic gradient descent steps, entire epochs, or any other sequence of operations.
@@
Online HPO.
However, perhaps one main advantage of explicitly taking an iterative view is that it enable recasting the HPO problem as an instance of online optimization [shalev2012online], albeit a non-standard one. Online HPO techniques attempt to speed up the HPO procedure by finding a performing hyperparamter schedule in one go, during a single execution of the learning algorithm. A key difficulty is that one does not have access to the true objective $\mathscr{M}$. Indeed, being able to compute (an unbiased estimate of) $\mathscr{M}$ requires terminating the learning algorithm, which would defeat the purpose of online HPO. In this sense, these techniques do not – or rather, cannot – strictly solve Problem (2.4)-(2.5), but rather rely on approximations, heuristics or surrogate objectives. This is akin to what we discussed in Chapter 5, however multi-fidelity techniques only use early observations to take decisions or fit models of the response surface, and do not change hyperparameters during execution. Examples of online HPO include [orabona2016coin, baydin_2017_online, luketina2016scalable, jaderberg2017population, lorenzo2017particle, mackay2019self, donini2019opt, vicol2021unbiased, micaelli2021gradient, chandra2022gradient], with gradient-based techniques, and especially forward iterative differentiation, being proposed by various authors.
Another stream of recent research focuses on tuning hyperparameters of online learning algorithms, such as contextual bandit [li2010contextual] methods like UCB [auer2002finite] or Thomson Sampling [agrawal2013thompson]. In these algorithms, hyperparameters affect the agent’s exploration-exploitation trade-off and must adapt in real time. bouneffouf2020hyper treat this as a nested bandit problem, tuning exploration parameters online via meta-bandits. ding2022syndicated generalize this idea with Syndicated Bandits, a multi-layer optimization scheme that tunes several hyperparameters simultaneously, with favorable regret guarantees. kang2024online further extend this line by framing tuning as a continuum-armed bandit problem, enabling efficient search in continuous spaces. These methods depart from gradient-based HPO discussed above by tightly coupling hyperparameter adaptation with decision-time feedback, offering principled and scalable alternatives tailored to the online nature of bandit learning.
Online HPO remains an area of active research.
@@ 8.3.1 Search Space
Instead of assigning operations to each node, liu2018darts assigns operations to every edge of the graph. Nodes that accumulate the feature vectors from all incoming edges [did not understand]. Their proposed method DARTS then uses gradient descent to select an operation for each edge.
@@ 8.3.2 Search strategies
In one of the first works on NAS, zoph2016neural used a controller based on a recurrent neural network to suggest new architectures. In each iteration, a set of candidate architectures is sampled and evaluated to collect observations, which are then used to update the controller using reinforcement learning. real2019regularized proposed an evolutionary approach that maintains a population of architectures (see Chapter 6). In each iteration, a random architecture is drawn from the population and mutated along a single dimension. While these approaches led to architectures outperforming state-of-the-art models at that point, the original methods required substantial computational resources since each network was trained from scratch.
@@ 8.4 Meta-learning
We discuss here meta-learning since, from a certain point of view, it shares some interesting similarities to HPO. This section only focuses on this point of view, in the case of supervised learning. For a more general overview of meta learning, see e.g., [hospedales:2021:metalearning_neural_networks] and references therein.
@@ 8.4.1 Definition
Meta-learning has its own specificity when compared to other settings dealing with several tasks. Compared to multi-task learning [Caruana94, Baxter2000, ruder:2017:overview_multitask_learning], the set of tasks is not fixed in advance. Compared to transfer learning or domain adaptation [pan2009survey, torrey2010transfer, kouw2018introduction], there are several “source” tasks. Typically, we are interested in the scenario where a very small dataset $\mathcal{D}\leavevmode\nobreak\ \sim p$ will be available for the new task (few-shot learning) or even a single example per class (one-shot learning). The main insight is that when the tasks are related, as captured by the distribution $\mathcal{P}$, meta-learning reduces the sample complexity of learning new tasks drawn by the same distribution [Baxter2000, maurer2016benefit]. Note that while HPO tipycally involves the optimization of a “black-box” function, both the meta-training and the meta-test phases in meta-learning more usually involve the optimization of functions with a know structure, enabling the application of gradient-based techniques, as discussed below.
8.4.2 Meta-learning of common representations
One effective approach to meta-learning consists of using past tasks to build strong representations, which can be later used for few-shot learning on the new task. We briefly discuss this approach here, as it is the most closely related to the HPO problem [franceschi2018bilevel]. As in [Baxter2000], let us introduce the family of hypothesis spaces $\mathbb{H}^{j}=\{\mathcal{H}^{j}\}$ where each $\mathcal{H}^{j}\in\mathbb{H}^{j}$ is a set of functions $h^{j}_{w,\lambda}:\mathcal{X}\mapsto\mathcal{Y}^{j}$ of the form $h^{j}_{w,\lambda}=\gamma_{w}^{j}\circ\Phi_{\lambda}$, i.e., obtained by composing a shared feature learning function $\Phi_{\lambda}:\mathcal{X}\mapsto\mathcal{F}$, parameterized by $\lambda$ (also called meta-parameter), with a task-specific function $\gamma_{w}^{j}:\mathcal{F}\mapsto\mathcal{Y}_{j}$, parameterized by $w$. The set $\mathcal{F}$ is a suitably chosen feature space, for example $\mathcal{F}=\mathbb{R}^{d}$. This type of function composition is common in several deep learning architectures where $\Phi$ is in charge of representation learning while $\gamma$ could be a classification or regression head. The multi-task network introduced in [Caruana94] and the feature learning neural network described in [Baxter2000] are structured in that way, although all the weights $\lambda$ and $w$ are optimized jointly, unlike the problem formulated below. Note that we have deliberately used the notation $\lambda$ for the parameters shared across the tasks, to highlight the fact that they play a role similar to hyperparameters. In order to construct a “response function,” for every source task $\mathcal{T}^{j}$, $j=1,\dots N$, we split the given dataset $\mathcal{D}^{j}$ into a training set $\mathcal{D}_{\text{tr}}^{j}$ and a validation set $\mathcal{D}_{\text{val}}^{j}$ and define